# A Gaussian process framework for testing general relativity with gravitational waves

### v1

@@ 3.3 Tests with other signals
4 Results from GWTC-3
We next apply our Gaussian process framework to the 60 binary black hole gravitational-wave events from GWTC-3 (Abbott et al., 2023) detected by both the H1 and L1 observatories. We perform parameter estimation to obtain posteriors on the binary parameters of the IMRPhenomPv2 waveform for 2 s of data around each event, using the same approach as in subsections 3.2 and 3.3. We produce residual data for each event by subtracting the maximum-likelihood estimate IMRPhenomPv2 waveform, and center the data at the maximum-likelihood merger time of each event. We perform Gaussian process regression on each of these data segments, as in the previous sections. We use the same priors on the Gaussian process hyper-parameters as outlined in subsection 3.1.
@@ 4 Results from GWTC-3
Figure 7: Posterior distribution of the Gaussian process hyper-parameters for GW150914 residual data. The contours are the $1$,$2$ and $3\sigma$ intervals. The posterior on $k_{0}$ does not exclude the lower bound of $10^{-46}$, indicating no signal is present in the data.
@@
    For each sample, we calculate the quadrature sum of the maximum $+$ and $\times$ polarisation strain amplitudes for the IMRPhenomPv2 waveform, $\delta s_{GR,\text{max}}$.
@@
We detail this process in Appendix C. We demonstrate this for the event with the most significant deviation only, GW190916_200658. We reweight 1500 randomly chosen hyper-posterior samples of $\Lambda$ to account for marginalising over the uncertainty in our GR waveform model (which is itself sampled by another set of 1000 posterior draws). This shifts the significance from $\ln{\mathcal{B}}=4.71$ to $\ln{\mathcal{B}}=3.85$. As this change is relatively minor, and since marginalisation over the uncertainty in the GR template tends to decrease the significance of potential deviations from GR, we do not perform this marginalisation for the other events studied.
5 Discussion
@@ Acknowledgements
This is LIGO document #P2500394. We acknowledge support from the Australian Research Council (ARC) Centres of Excellence CE170100004 and CE230100016, as well as ARC LE210100002, and ARC DP230103088. L.P. receives support from the Australian Government Research Training Program. S.S is a recipient of an ARC Discovery Early Career Research Award (DE220100241). This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459.
@@
The data underyling this article are publicly available at https://www.gw-openscience.org.
@@ Appendix B Reconstructing the deviation
We wish to reconstruct the deviation from GR $\delta s$ given residual data $\delta h$ – that is, we want to construct ${p}(\delta s|\delta h)$. We first expand Eq. 14 using Eqs. 15 and 2.3, giving
${\cal L}(\delta h|\delta s)=\frac{1}{2\pi\text{det}(\textbf{{N}})}\exp\Big{(}-(\delta h-\delta s)^{\dagger}\textbf{{N}}^{-1}(\delta h-\delta s)\Big{)}\,\frac{1}{2\pi\text{det}(\textbf{{S}}(\Lambda))}\exp\Big{(}-\delta s^{\dagger}\textbf{{S}}^{-1}(\Lambda)\delta s\Big{)}.$ (B1)
@@
[…]
@@
[…]
@@
$p(\delta s|\delta h)\propto$ $\sum_{k}\frac{1}{2\pi\text{det}(\textbf{{S}}(\Lambda_{k}))}\frac{1}{2\pi\text{det}(\textbf{{N}})}\frac{1}{{\cal L}(\Lambda_{k}|\delta h)}\exp\Big{(}-(\delta h-\delta s)^{\dagger}\textbf{{N}}^{-1}(\delta h-\delta s)\Big{)}\exp\Big{(}-\delta s^{\dagger}\textbf{{S}}^{-1}(\Lambda_{k})\delta s\Big{)}.$ (B8)
@@
×$\quad\times\exp\Big{(}-\frac{1}{2}\delta h^{\dagger}\textbf{{N}}^{-1}(\textbf{{N}}^{-1}\textbf{{S}}^{-1}(\Lambda_{k}))^{-1}\textbf{{N}}^{-1}-I)\delta h\Big{)},$ (B9)
@@
$\Sigma=\textbf{{N}}^{-1}+\textbf{{S}}^{-1}(\Lambda_{k}).$ (B12)
@@ Appendix C Marginalising over uncertainty in the gravitational waveform
[…]
@@
where we have canceled the factor $\pi(\theta_{i})$ in the numerator and denominator. Here, $\mathcal{Z}_{\theta}$ is the evidence for the gravitational waveform model. Since we desire the reweighted evidence $\mathcal{Z}_{\text{RW,GP}}$ that takes into account waveform uncertainty:
[…]
@@
[…]
@@
Subsituting into the equation for $\mathcal{B}_{\text{RW}}$ and simplifying gives
@@
For GW190916_200658, the number of hyper-posterior samples is $1500$, and $n_{\text{eff}}=356$, indicating a reweighting efficiency of $24\%$, which we believe is sufficient for this demonstration.

### v2

@@ III.3 Tests with other signals
IV Results from GWTC-3
We next apply our Gaussian process framework to the 60 binary black hole gravitational-wave events from GWTC-3 (Abbott et al., 2023) detected by both the H1 and L1 observatories. We do not consider any glitch-subtracted data frames, instead using the publicly available GWOSC data (Abbott and others, 2023). We perform parameter estimation to obtain posteriors on the binary parameters of the IMRPhenomPv2 waveform for 2 s of data around each event, using the same approach as in subsections III.2 and III.3. We produce residual data for each event by subtracting the maximum-likelihood estimate IMRPhenomPv2 waveform, and center the data at the maximum-likelihood merger time of each event. We perform Gaussian process regression on each of these data segments, as in the previous sections. We use the same priors on the Gaussian process hyper-parameters as outlined in subsection III.1.
@@ IV Results from GWTC-3
Figure 7: Posterior distribution of the Gaussian process hyper-parameters for GW150914 residual data. The contours are the 11,22 and $3\sigma$ intervals. The posterior on $k_{0}$ does not exclude the lower bound of $10^{-46}$, indicating no signal is present in the data.
@@
    For each sample, we calculate the quadrature sum of the maximum ++ and $\times$ polarisation strain amplitudes for the IMRPhenomPv2 waveform, $\delta s_{GR,\text{max}}$.
@@
We detail this process in Appendix C. We demonstrate this for the event with the most significant deviation only, GW190916_200658, and find the significance shifts from $\ln{\mathcal{B}}=4.71$ to $\ln{\mathcal{B}}=3.85$. As this change is relatively minor, and since marginalisation over the uncertainty in the GR template tends to decrease the significance of potential deviations from GR, we do not perform this marginalisation for the other events studied.
V Discussion
@@
Similar to Ashton (2023), it may be possible to extend this framework to model noise glitches in gravitational-wave data. This would require revision of the kernel used in analysis, to either loosen the coherency constraint between detectors, or introduce new kernel functions/hyperparameters to more suitably model typical glitches. For example, one could allow for additional characteristic frequencies, allow for time-dependent (‘chirping’) characteristic frequencies, or allow the peak of the signal to be different to the merger time. We leave this as an avenue for future research.
@@ Acknowledgements
This is LIGO document #P2500394. We acknowledge support from the Australian Research Council (ARC) Centres of Excellence CE170100004 and CE230100016, as well as ARC LE210100002, and ARC DP230103088. L.P. and S.Y.C. receive support from the Australian Government Research Training Program. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459.
@@
The data underlying this article are publicly available at https://www.gw-openscience.org.
@@ Appendix B Reconstructing the deviation
We wish to reconstruct the deviation from GR $\delta s$ given residual data $\delta h$ – that is, we want to construct ${p}(\delta s|\delta h)$. We first expand Eq. 14 using Eqs. 15 and II.3, giving
${\cal L}(\delta h|\delta s)\propto\frac{1}{\text{det}(\textbf{{N}})}\exp\Big(-(\delta h-\delta s)^{\dagger}\textbf{{N}}^{-1}(\delta h-\delta s)\Big)\,\frac{1}{\text{det}(\textbf{{S}}(\Lambda))}\exp\Big(-\delta s^{\dagger}\textbf{{S}}^{-1}(\Lambda)\delta s\Big).$ (B1)
@@
${p}(\delta s|\delta h)=$ $\int d\Lambda\,p(\delta s|\delta h,\Lambda)$ (B3)
$=$ $\int d\Lambda\,\frac{{\cal L}(\delta h|\delta s)\pi(\delta s|\Lambda)\pi(\Lambda)}{\cal Z}.$ (B4)
@@
$p(\delta s|\delta h)=$ $\int d\Lambda\,\frac{{\cal L}(\delta h|\delta s)\pi(\delta s|\Lambda)\pi(\Lambda)}{\cal Z}\frac{{p}(\Lambda|\delta h)}{{p}(\Lambda|\delta h)},$ (B5)
@@
$p(\delta s|\delta h)\propto$ $\sum_{k}\frac{1}{\text{det}(\textbf{{S}}(\Lambda_{k}))}\frac{1}{\text{det}(\textbf{{N}})}\frac{1}{{\cal L}(\Lambda_{k}|\delta h)}\exp\Big(-(\delta h-\delta s)^{\dagger}\textbf{{N}}^{-1}(\delta h-\delta s)\Big)\exp\Big(-\delta s^{\dagger}\textbf{{S}}^{-1}(\Lambda_{k})\delta s\Big).$ (B8)
@@
$\quad\times\exp\Big(-\frac{1}{2}\delta h^{\dagger}\textbf{{N}}^{-1}(\textbf{{N}}^{-1}\textbf{{S}}^{-1}(\Lambda_{k}))^{-1}\textbf{{N}}^{-1}-I)\delta h\Big),$ (B9)
@@
$\Sigma=(\textbf{{N}}^{-1}+\textbf{{S}}^{-1}(\Lambda_{k}))^{-1}.$ (B12)
@@ Appendix C Marginalising over uncertainty in the gravitational waveform
${\cal L}(d|\Lambda)=\int d\theta{\cal L}(d|\Lambda,\theta)\pi(\theta)\frac{p(\theta|d)}{p(\theta|d)}.$ (C2)
@@
where we have canceled the factor $\pi(\theta_{i})$ in the numerator and denominator. Here, $\mathcal{Z}_{\theta}$ is the evidence for the gravitational waveform model. We desire the reweighted evidence $\mathcal{Z}_{\text{RW,GP}}$ that takes into account waveform uncertainty:
$\mathcal{Z}_{\text{RW,GP}}=\int{d\Lambda{\cal L}(d|\Lambda)\pi(\Lambda)}\approx\int d\Lambda\frac{\mathcal{Z}_{\theta}}{N_{\theta}}\sum_{i=1}^{N_{\theta}}\frac{{\cal L}(d|\Lambda,\theta_{i})}{{\cal L}(d|\theta_{i})}\pi(\Lambda).$ (C5)
@@
$\mathcal{Z}_{\text{RW,GP}}\approx\frac{\mathcal{Z}_{\theta}}{N_{\theta}}\int d\Lambda\sum_{i=1}^{N_{\theta}}\frac{{\cal L}(d|\Lambda,\theta_{i})}{{\cal L}(d|\theta_{i})}\frac{p(\Lambda|d,\theta_{\text{ML}})}{p(\Lambda|d,\theta_{\text{ML}})}\pi(\Lambda).$ (C6)
@@
Substituting into the equation for $\mathcal{B}_{\text{RW}}$ and simplifying gives
@@
We now demonstrate this for the event with the highest significance, GW190916_200658. We reweight 1500 randomly chosen hyper-posterior samples of $\Lambda$ to account for marginalising over the uncertainty in our GR waveform model (which is itself sampled by another set of 1000 posterior draws). We show the effects of this on both the hyper-parameter posteriors and the reconstructed deviation in Figures 10 and 11. We find the reconstructed deviation broadens minimally, and its significance shifts from $\ln{\mathcal{B}}=4.71$ to $\ln{\mathcal{B}}=3.85$, which we believe is relatively minor, and thus do not perform this treatment for other events. The number of hyper-posterior samples is $1500$, and $n_{\text{eff}}=356$, indicating a reweighting efficiency of $24\%$, which we believe is sufficient for this demonstration.
Figure 10: Posterior distribution of the Gaussian process hyper-parameters for GW190916_200658 residual data. The contours are the 11,22 and $3\sigma$ intervals. Shown in black are the posteriors for the initial max-likelihood estimate run for the GR waveform parameters $\theta$, and shown in orange are the re-weighted posteriors, taking into account the uncertainty in $\theta$.
Figure 11: Reconstruction of the whitened deviation strain $\delta{s}$ for GW190916_200658 residual data. The blue curve is the whitened residual strain in each detector. The black curve is the median estimate of $\delta{s}$ using Gaussian process regression for the max-likelihood estimate of the GR waveform parameters $\theta$, and the grey shaded region is the $90\%$ credible interval on $\delta{s}$ for this estimate. The orange shaded region is the $90\%$ credible interval on $\delta{s}$ marginalised over the uncertainty in $\theta$.